OpenStudy (anonymous):

What is the improper fraction that falls between 22/5 and 9/2

OpenStudy (anonymous):

9/2>x>22/5 18/10>x>44/10 Take your pic

OpenStudy (anonymous):

pick*

sammixboo (sammixboo):

Okay lets convert the fractions \(\sf \large \color{magenta}{\frac{22}{5}}\) and \(\sf \large \color{magenta}{\frac{9}{2}}\) into mix fractions... \(\sf \large \color{magenta}{\frac{22}{5}}\) ---- \(\sf \large \color{magenta}{4\frac{2}{5}}\) \(\sf \large \color{magenta}{\frac{9}{2}}\) ---- \(\sf \large \color{magenta}{4\frac{1}{2}}\) Now we need to make these numbers have the common denominator of 20 \(\sf \large \color{magenta}{4\frac{1 x 10}{2 x 10}}\) ---- \(\sf \large \color{magenta}{4\frac{10}{20}}\) \(\sf \large \color{magenta}{4\frac{2 x 4}{5 x 4}}\) ---- \(\sf \large \color{magenta}{4\frac{8}{20}}\)

sammixboo (sammixboo):

Now what mixed fraction in between \(\sf \large \color{magenta}{4\frac{8}{20}}\) and \(\sf \large \color{magenta}{4\frac{10}{20}}\)?